three years ago, the sum of the ages of marlon and carol was 26. the difference between their ages 2 years from now will be 4. find marlons and carols present ages. assume that marlon is older than carol.
Answer by ankor@dixie-net.com(22740)

let m = M's present age
let c = C's present age
:
Write an equation and simplify it, for each statement.
:
"three years ago, the sum of the ages of marlon and carol was 26."
(m-3) + (c-3) = 26
m + c - 6 = 26
m + c = 26 + 6
m + c = 32
:
"the difference between their ages 2 years from now will be 4."
(m+2) - (c+2) = 4
m - c + 2 - 2 = 4
m - c = 4
:
Add the two equation
m - c = 4
m + c = 32
------------adding eliminates c, find m
2m = 36
m = 36/2
m = 18 yrs is M's age
:
I'll let you find C's age; check the solutions in both statements